dakchicken:

A quadratic function f(x) has two zeros at (0, 0) and (4, 0) . The extreme value of f(x) is a maximum at (2, 12). Which function among the following fits all the properties above?

snowflake0531:

which are teh choices?

dakchicken:

f (x) = −3x (x − 4) f (x) = x (x + 4) f (x) = 3x (x − 4) f (x) = −x (x + 4)

jhonyy9:

so like a first step check what function has zeros of (0,0) and (4,0) and the maxim at (2,12) hope help you

snowflake0531:

Firslty, you should know that this will be a negative thing, because it has to go through the x axis, but also a maximum point at (2,12)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step check what function has zeros of (0,0) and (4,0) and the maxim at (2,12) hope help you \(\color{#0cbb34}{\text{End of Quote}}\) @AZ opinion pls - ty

cashmoneyAX:

wsp yall

AZ:

I agree. Also you should know that for it to have a maximum, a < 0

1 attachment
dakchicken:

so whts the answer

AZ:

Here are your answer choices f (x) = −3x (x − 4) f (x) = x (x + 4) f (x) = 3x (x − 4) f (x) = −x (x + 4) Can you plug in (4, 0) into each one and tell me for which answer choices that it makes it work out?

AZ:

To show you what I mean, I'll do answer choice A f(x) = -3x (x - 4) Plug in (4, 0) We get 0 = -3* (4) * (4 - 4) 0 =-3*4* 0 0 = 0 So answer choice A works what other answer choices work and which ones will we have to eliminate?

dakchicken:

thank you

AZ:

we haven't completed it yet and I didn't say that was the answer there's one more answer choice that has a solution for (4, 0) we then need to figure out which one has the maximum at (2, 12)